An aluminum sphere is 8.15 cm in diameter Part A What will be its change in volume if it is heated from 40 C to 150 C? Express your answer using two significant figures. Submit Previosis Answers Request Answer
Answer #1

volume of the sphere, V (4/3)*pi*(d/2)^3

= (4/3)*pi*(8.15/2)^3

= 283 cm^3

coefficient of volume expansion of aluminum, gamma = 66*10^-6 K^-1

using the equation,

delta_V = V*gamma*(T2 - T1)

= 283*66*10^-6*(150 - 40)

= 2.05 cm^3 <<<<<<<<<<-------------Answer